OpenStudy (anonymous):

Blake has a total of 64 DVD's and CD's. If the number of CD's is eight more than three times the number of DVD's, how many DVD's does he have?

OpenStudy (anonymous):

Let x be the number of CDs and y the number of DVDs. Then,\[x+y=64\]\[x=3y+8\] \[x=64-y\]\[64-y=3y+8\] \[x=64-y\]\[4y=56\] \[x=50\]\[y=14\] So Blake has 14 DVDs.